# Generalizing Consistency Policy to Visual RL with Prioritized Proximal Experience Regularization

### v1

@@ Abstract
With high-dimensional state spaces, visual reinforcement learning (RL) faces significant challenges in exploitation and exploration, resulting in low sample efficiency and training stability. As a time-efficient diffusion model, although consistency models have been validated in online state-based RL, it is still an open question whether it can be extended to visual RL. In this paper, we investigate the impact of non-stationary distribution and the actor-critic framework on consistency policy in online RL, and find that consistency policy was unstable during the training, especially in visual RL with the high-dimensional state space. To this end, we suggest sample-based entropy regularization to stabilize the policy training, and propose a consistency policy with prioritized proximal experience regularization (CP3ER) to improve sample efficiency. CP3ER achieves new state-of-the-art (SOTA) performance in 21 tasks across DeepMind control suite and Meta-world. To our knowledge, CP3ER is the first method to apply diffusion/consistency models to visual RL and demonstrates the potential of consistency models in visual RL. More visualization results are available at https://jzndd.github.io/CP3ER-Page/.
@@ 1 Introduction
As an emerging generation model, the diffusion model [11] stands out in fields such as image generation [12, 13, 14] and video generation [15, 16] with its ability to model complex distributions and ease of sampling characteristics. These properties have also been explored for learning a complex policy [17]. For example, diffusion models are used to imitation learning to learn the diverse expert policies [18, 19] or trajectories [20, 21, 22, 23] in datasets. In addition, due to their excellent expressive and data generation abilities, diffusion models are often employed to address policy constraints [24, 25, 26] and data scarcity [21, 27, 28] in offline RL. Most of these applications are limited to offline learning due to the demand for pre-collected datasets to train diffusion models.
@@
In this paper, we investigate the impact of non-stationary dataset and the actor-critic framework on consistency policy. By analyzing the dormant rate [29] of the policy network, we find that the non-stationary of training data is not the main factor affecting the instability of consistency policy, while the Q-loss in the actor-critic framework leads to a sharp increase in the dormant rate of the policy network, resulting in the loss of complex expression ability, which is particularly significant in visual RL tasks. To address the above issues, we suggest sample-based entropy regularization to stabilize the policy training and propose the prioritized proximal experience regularization, which uses weighted sampling to construct an appropriate proxy policy for policy regularization and achieves sample-efficiency consistency policy. Overall, our contributions are as follows:
@@ 2.1 Diffusion Model in Reinforcement Learning
Due to its high-quality sample generation ability and training stability, diffusion models [11] have been widely applied in fields such as image generation, video generation, and natural language processing and have also been promoted in RL. Since the diffusion model can represent complex distribution in datasets, it is commonly used in offline RL to model behavior policies [18, 25, 35] or expected policies [34, 36, 37, 38] to meet the requirements of diversity policy constraints and achieve a balance between constraint and exploitation. The diffusion model can also model trajectory distribution [20, 39], achieving specified trajectory generation under different guidance. In addition, diffusion models are also employed to generate data to augment limited training data [27, 28].
@@ 3.3 Dormant Ratio of Neural Networks
The expressive ability is crucial for training the policy with RL. [29] proposes the concept of dormant rate , which quantifies the expression ability of a neural network by calculating the proportion of dormant neurons in the neural networks.
@@
Here is the activation function of the -th neuron in the -th layer. is the distribution of the input . In the following sections of this paper, we use the dormant rate to evaluate the expression ability of consistency policy during the training.
@@ 4 Is Consistency-AC Applicable to Visual RL?
Does the non-stationary distribution in online RL affect the training of consistency models? Unlike offline RL, online RL does not have pre-collected datasets. The data distribution for training the policy is constantly changing with policy improvement. So, whether this non-stationarity distribution affects the training of consistency models is a question that needs to be explored. In order to investigate the impact of non-stationarity of data for consistency model training, we record the dormant rate of the policy network during consistency model training under two different settings: online training and offline training. We employ two tasks (MuJoCo Halfcheetah and MuJoCo Walker2d) and conduct 4 random seeds for each setting. In order to eliminate the impact of Q-loss, we follow the behavior clone setting and train the consistency model with consistency loss [32] using data from offline datasets or online replay buffers. The distribution of the data in the replay buffer varies with policy improvement. The results are shown in Figure 1. Although there is a difference in the dormant rates between online and offline learning settings in the Halfcheetah task, the overall trend is the same. We speculate that this difference is caused by the diversity of the samples included in the dataset. For the Walker2d task, the dormant rates are nearly the same under two different settings. Therefore, we can infer that the non-stationary distribution of online RL does not significantly affect the consistency model training.
@@
Is the actor-critic framework suitable for training consistency policy? The actor-critic framework is a highly effective policy training framework for online RL, in which the policy network achieves policy improvement by maximizing the value function. Some works [33, 34] directly apply consistency models to this framework. Although they achieve good results in RL tasks with low dimensional state spaces, whether the actor-critic training framework is compatible with consistency model training remains a question that needs further investigation. To evaluate the impact of the actor-critic framework on the training of consistency models, we compare the dormant rates of policy networks under the consistency loss and Q-loss settings under the actor-critic framework. The results are shown in Figure LABEL:fig:online_state_dormant. When training the policy with the consistency loss, the dormant rates of the network show a trend of first decreasing and then increasing. This means that the policy learns the distribution from the data and then overfits the distribution. When training the policy with the Q-loss, the dormant rate of the policy network will rapidly increase and maintain a high value, which means that the policy network will quickly fall into local optima, making the policy no longer change. In addition, we can also see that when using the Q-loss, the variance of the dormant rates is relatively large under different random seeds. When the dormant rate is low, the policy network can iterate properly to learn good policy. Therefore, we can determine that the Q-loss under the actor-critic framework will destabilize the consistency policy training.
@@
Will high-dimensional state space exacerbate the degradation phenomenon of consistency policy? Compared to RL with low dimensional state space, training stability in visual RL is still a challenge. In order to investigate whether the degradation phenomenon of consistency policy will become more significant under visual RL tasks, we compare the dormant rates of the policy networks with the state as input and image as input on 2 tasks (Walker-walk and Cheetah-run in DeepMind control suite) under the setting of online learning. During the training process, only the Q-loss was used. To maintain consistency in the settings, we only count the dormant rate of the multilayer perceptron (MLP) of the policy newtork in the image-based settings. The results are shown in Figure LABEL:fig:online_image_dormant. Similar to using the state as input, in visual RL with the image as input, most of the neurons in the MLP of the policy network go dormant. Unlike the high variance of the former, the dormant rates of consistency policy network in visual RL maintain a low variance and a high value. This indicates that there have been almost no successful trials under different random seeds. Therefore, we can infer that visual RL will exacerbate the instability of consistency policy training caused by the Q-loss under the actor-critic framework.
@@ 5 Consistency Policy with Prioritized Proximal Experience Regularization
Consistency Policy with Entropy Regularization To solve the problem of consistency policy quickly falling into local optima under the influence of the Q-loss, we introduce policy regularization to stabilize policy improvement. Here, we employ entropy regularization to constrain policy behavior. The objective of RL is:
@@
Prioritized Proximal Experience Regularization When the proxy distribution is uniform, this method approximates the maximum entropy consistency policy (MaxEnt CP). It should be noted that the difference between the proxy distribution and the optimal policy distribution can lead to difficulty in optimizing the above objectives. When the proxy distribution is far from the optimal policy or the proxy distribution is complex, the above optimization objectives require more samples to converge to better results. To better balance training stability and sample efficiency, we propose the prioritized proximal experience regularization (PPER). Specifically, when sampling data from the replay buffer, we design sampling weight for each data instead of sampling the data uniformly.
@@
where is the estimated value distribution. According to the equation (9), we need to sample target Q-values and update the value distribution. Different from [53], we sample only one next action instead of multiple actions to reduce the time cost and find that this simplification can achieve good experiment results. Considering the simplicity and efficiency of DrQ-v2 [54], our proposed consistency policy with prioritized proximal experience regularization (CP3ER) is built based on DrQ-v2. The framework is shown as Figure 3 (a). We sample the data from the replay buffer, and augment the image with the techniques in DrQ-v2. Thanks to the natural randomness of the action from consistency policy, our method no longer requires additional exploration strategies. In addition, we only used a single Q-network to estimate the mean and variance of the mixture of Gaussian instead of double Q-network. We consider prioritized proximal experience regularization when updating consistency policy and used equation (9) when training the Q-network, which differs from DrQ-v2. The complete algorithm is included in the appendix A.1.
@@ 6.1 Visual Continuous Control Tasks
Environment Setup We evaluate the methods on 21 visual control tasks from DeepMind control suite [55] and Meta-world [56]. We split these tasks into three domains, including 8 medium-level tasks in the DeepMind control suite, 7 hard-level tasks in the DeepMind control suite, and 6 tasks in the Meta-world. The details of each domain are included in the appendix B.
Baselines We compare current advanced model-free visual RL methods, including DrQ-v2 [54], ALIX [57] and TACO [58]. The more detailed results are shown in the appendix B.
@@ 6.1.1 Does CP3ER have performance advantages compared to current SOTA methods?
Medium-level tasks in DeepMind control suite. We evaluated CP3ER on 8 medium-level tasks [54] in DeepMind control suite. The results are shown in Figure 4. From the left part of Figure 4, it can be seen that compared to the current SOTA methods, our proposed CP3ER has achieved better sample efficiency. It should be noted that TACO uses auxiliary losses of action and state representation during training to improve sample efficiency. Moreover, our proposed CP3ER uses no additional losses or exploration strategies. On the right part of Figure 4, we compare the mean, Interquartile Mean (IQM), median, and optimal gap of these methods. CP3ER has significant advantages in all metrics and has more minor variance. This means that CP3ER has better training stability.
Hard-level tasks in DeepMind control suite. We also evaluate CP3ER on seven challenging tasks [52, 54] in the DeepMind control suite. It should be noted that all tasks here only train 5M frames, rather than the commonly used 30M frames in other work[52, 54]. This means it is a very hard challenge. From the results on the left part of Figure 5, it can be seen that most methods have yet to learn effective policy within 5M frames. Our proposed CP3ER surpasses the performance of all methods without relying on any additional loss or exploration strategies. Moreover, it has significant advantages on all metrics including mean, IQM, median, and optimal gap.
@@ 6.2.1 Can policy regularization improve the behavior of the policy during training?
Action distribution analysis with toy example. In order to further explore the impact of policy regularization on the training, we borrow the 1D continuous bandit problem [9] to analyze the policy network behavior. The green curve in Figure 7a (a) shows the reward function. Within a narrow range of actions, the agent receives higher rewards, while within a broader range, the agent can only receive suboptimal rewards. Therefore, the policy needs strong exploration ability to achieve the highest return. We compare Gaussian policy with entropy regularization (MaxEnt GP), Consistency-AC[33] and consistency policy with entropy regularization (MaxEnt CP), and record the action distribution during the training. As shown in Figure 7a, Consistency-AC quickly converges to the local optimal value with the Q-loss. Policy regularization ensures the diversity of action distribution during consistency policy training, preventing the policy from falling into local optima too early. Moreover, consistency policy has robust exploration compared to the Gaussian policy and achieves higher returns.
Dormant rate analysis. We have shown that the Q-loss can rapidly increase the dormant rate of the consistency policy network, leading to a loss of policy diversity. In order to analyze whether entropy regularization can alleviate the phenomenon, we record the dormant rates of the policy networks during the training in 3 tasks. The results are shown in Figure 8. Compared to the rapid increase in the dormant rate in Consistency-AC, CP3ER has a lower dormant rate, which means that entropy regularization can effectively reduce the dormant rates of consistency policy.
@@ 7 Conclusion
In this paper, we analyze the problems faced by extending consistency policy to visual RL under the actor-critic framework and discover the phenomenon of the collapse of consistency policy during training under the actor-critic framework by analyzing the dormant rate of the neural networks. To address this issue, we propose a consistency policy with prioritized proximal experience regularization (CP3ER) that effectively alleviates the training collapse problem of consistency policy. The method is evaluated on 21 visual control tasks and shows significantly better sample efficiency and performance than the current SOTA methods. It is worth mentioning that, to our knowledge, our proposed CP3ER is the first method to apply diffusion/consistency models to visual RL tasks.
@@
Appendix A Implementation details

### v2

@@ Abstract
With high-dimensional state spaces, visual reinforcement learning (RL) faces significant challenges in exploitation and exploration, resulting in low sample efficiency and training stability. As a time-efficient diffusion model, although consistency models have been validated in online state-based RL, it is still an open question whether it can be extended to visual RL. In this paper, we investigate the impact of non-stationary distribution and the actor-critic framework on consistency policy in online RL, and find that consistency policy was unstable during the training, especially in visual RL with the high-dimensional state space. To this end, we suggest sample-based entropy regularization to stabilize the policy training, and propose a consistency policy with prioritized proximal experience regularization (CP3ER) to improve sample efficiency. CP3ER achieves new state-of-the-art (SOTA) performance in 21 tasks across DeepMind control suite and Meta-world. To the best of our knowledge, CP3ER is the first method to apply diffusion/consistency models to visual RL and demonstrates the potential of consistency models in visual RL. Our project page is hosted at https://jzndd.github.io/CP3ER-Page/.
@@ 1 Introduction
As an emerging generation model, the diffusion model [11] stands out in fields such as image generation [12, 13, 14] and video generation [15, 16] with its ability to model complex distributions and ease of sampling characteristics. These properties have also been explored for learning a complex policy [17]. For example, diffusion models are used to imitat e the diverse expert policies [18, 19] or trajectories [20, 21, 22, 23] in datasets. In addition, due to their excellent expressive and data generation abilities, diffusion models are often employed to address policy constraints [24, 25, 26] and data scarcity [21, 27, 28] in offline RL. Most of these applications are limited to offline learning due to the demand for pre-collected datasets to train diffusion models.
@@
In this paper, we investigate the impact of non-stationary dataset and the actor-critic framework on consistency policy. By analyzing the dormant ratio [29] of the policy network, we find that the non-stationary of training data is not the main factor affecting the instability of consistency policy, while the Q-loss in the actor-critic framework leads to a sharp increase in the dormant ratio of the policy network, resulting in the loss of complex expression ability, which is particularly significant in visual RL tasks. To address the above issues, we suggest sample-based entropy regularization to stabilize the policy training and propose the prioritized proximal experience regularization, which uses weighted sampling to construct an appropriate proxy policy for policy regularization and achieves sample-efficiency consistency policy. Overall, our contributions are as follows:
@@ 2.1 Diffusion Model in Reinforcement Learning
Due to its high-quality sample generation ability and training stability, diffusion models [11] have been widely applied in fields such as image generation, video generation, and natural language processing and have also been promoted in RL. Since the diffusion model can represent complex distribution in datasets, it is commonly used in offline RL to model behavior policies [18, 25, 35] or expected policies [34, 36, 37, 38] to meet the requirements of diversity policy constraints and achieve a balance between constraint and exploitation. The diffusion model can also model trajectory distribution [20, 39, 40], achieving specified trajectory generation under different guidance. In addition, diffusion models are also employed to generate data to augment limited training data [27, 28].
@@ 3.3 Dormant Ratio of Neural Networks
The expressive ability is crucial for training the policy with RL. [29] proposes the concept of dormant ratio , which quantifies the expression ability of a neural network by calculating the proportion of dormant neurons in the neural networks.
@@
Here is the activation function of the -th neuron in the -th layer. is the distribution of the input . In the following sections of this paper, we use the dormant ratio to evaluate the expression ability of consistency policy during the training.
As introduced in [29], the dormant ratio of a neural network indicates the proportion of inactive neurons and is typically used to measure the activity of the network. A higher dormant ratio implies fewer active neurons in the network, implying the network’s capacity and expressiveness are damaged. In RL, the episode return is closely related to the dormant ratio of the policy network. A higher dormant ratio results in more lazy action outputs, inactive agent behavior, and lower episode returns; conversely, when policy performance is good, the policy network is usually more active, and the dormant ratio is typically lower. This phenomenon has been reported in [29, 53, 54, 55].
@@ 4 Is Consistency-AC Applicable to Visual RL?
Does the non-stationary distribution in online RL affect the training of consistency models? Unlike offline RL, online RL does not have pre-collected datasets. The data distribution for training the policy is constantly changing with policy improvement. So, whether this non-stationarity distribution affects the training of consistency models is a question that needs to be explored. In order to investigate the impact of non-stationarity of data for consistency model training, we record the dormant ratio of the policy network during consistency model training under two different settings: online training and offline training. We employ two tasks (MuJoCo Halfcheetah and MuJoCo Walker2d) and conduct 4 random seeds for each setting. In order to eliminate the impact of Q-loss, we follow the behavior clone setting and train the consistency model with consistency loss [32] using data from offline datasets or online replay buffers. The distribution of the data in the replay buffer varies with policy improvement. The results are shown in Figure 1. Although there is a difference in the dormant ratios between online and offline learning settings in the Halfcheetah task, the overall trend is the same. We speculate that this difference is caused by the diversity of the samples included in the dataset. For the Walker2d task, the dormant ratios are nearly the same under two different settings. Therefore, we can infer that the non-stationary distribution of online RL does not significantly affect the consistency model training.
@@
Is the actor-critic framework suitable for training consistency policy? The actor-critic framework is a highly effective policy training framework for online RL, in which the policy network achieves policy improvement by maximizing the value function. Some works [33, 34] directly apply consistency models to this framework. Although they achieve good results in RL tasks with low dimensional state spaces, whether the actor-critic training framework is compatible with consistency model training remains a question that needs further investigation. To evaluate the impact of the actor-critic framework on the training of consistency models, we compare the dormant ratios of policy networks under the consistency loss and Q-loss settings under the actor-critic framework. The results are shown in (a) and (b) of Figure 2, the solid line shows the dormant ratio of the policy while the dashed line shows the performance of the policy. When training the policy with the consistency loss, the dormant ratios of the network show a trend of first decreasing and then increasing. This means that the policy learns the distribution from the data and then overfits the distribution. When training the policy with the Q-loss, the dormant ratio of the policy network will rapidly increase and maintain a high value, which means that the policy network will quickly fall into local optima, making the policy no longer change. In addition, we can also see that when using the Q-loss, the variance of the dormant ratios is relatively large under different random seeds. When the dormant ratio is low, the policy network can iterate properly to learn good policy. Therefore, we can determine that the Q-loss under the actor-critic framework will destabilize the consistency policy training.
@@
Will high-dimensional state space exacerbate the degradation phenomenon of consistency policy? Compared to RL with low dimensional state space, training stability in visual RL is still a challenge. In order to investigate whether the degradation phenomenon of consistency policy will become more significant under visual RL tasks, we compare the dormant ratios of the policy networks with the state as input and image as input on 2 tasks (Walker-walk and Cheetah-run in DeepMind control suite) under the setting of online learning. During the training process, only the Q-loss was used. To maintain consistency in the settings, we only count the dormant ratio of the multilayer perceptron (MLP) of the policy newtork in the image-based settings. The results are shown in (c) and (d) of Figure 2. Similar to using the state as input, in visual RL with the image as input, most of the neurons in the MLP of the policy network go dormant. Unlike the high variance of the former, the dormant ratios of consistency policy network in visual RL maintain a low variance and a high value. This indicates that there have been almost no successful trials under different random seeds. Therefore, we can infer that visual RL will exacerbate the instability of consistency policy training caused by the Q-loss under the actor-critic framework.
@@ 5 Consistency Policy with Prioritized Proximal Experience Regularization
Consistency Policy with Entropy Regularization. To solve the problem of consistency policy quickly falling into local optima under the influence of the Q-loss, we introduce policy regularization to stabilize policy improvement. Here, we employ entropy regularization to constrain policy behavior. The objective of RL is:
@@
Prioritized Proximal Experience Regularization. When the proxy distribution is uniform, this method approximates the maximum entropy consistency policy (MaxEnt CP). It should be noted that the difference between the proxy distribution and the optimal policy distribution can lead to difficulty in optimizing the above objectives. When the proxy distribution is far from the optimal policy or the proxy distribution is complex, the above optimization objectives require more samples to converge to better results. To better balance training stability and sample efficiency, we propose the prioritized proximal experience regularization (PPER). Specifically, when sampling data from the replay buffer, we design sampling weight for each data instead of sampling the data uniformly.
@@
where is the estimated value distribution. According to the equation (9), we need to sample target Q-values and update the value distribution. Different from [56], we sample only one next action instead of multiple actions to reduce the time cost and find that this simplification can achieve good experiment results. Considering the simplicity and efficiency of DrQ-v2 [57], our proposed consistency policy with prioritized proximal experience regularization (CP3ER) is built based on DrQ-v2. The framework is shown as Figure 3 (a). We sample the data from the replay buffer, and augment the image with the techniques in DrQ-v2. Thanks to the natural randomness of the action from consistency policy, our method no longer requires additional exploration strategies. In addition, we only used a single Q-network to estimate the mean and variance of the mixture of Gaussian instead of double Q-network. We consider prioritized proximal experience regularization when updating consistency policy and used equation (LABEL:mog_q_loss) when training the Q-network, which differs from DrQ-v2. The complete algorithm is included in the appendix B.1.
@@ 6.1 Visual Continuous Control Tasks
Environment Setup. We evaluate the methods on 21 visual control tasks from DeepMind control suite [58] and Meta-world [59]. We split these tasks into three domains, including 8 medium-level tasks in the DeepMind control suite, 7 hard-level tasks in the DeepMind control suite, and 6 tasks in the Meta-world. The details of each domain are included in the appendix C.
Baselines. We compare current advanced model-free visual RL methods, including DrQ-v2 [57], ALIX [60] and TACO [61]. The more detailed results are shown in the appendix C.
@@ 6.1.1 Does CP3ER have performance advantages compared to current SOTA methods?
Medium-level tasks in DeepMind control suite. We evaluate CP3ER on 8 medium-level tasks [57] in DeepMind control suite. The results are shown in Figure 4. From the left part of Figure 4, it can be seen that compared to the current SOTA methods, our proposed CP3ER has achieved better sample efficiency. It should be noted that TACO uses auxiliary losses of action and state representation during training to improve sample efficiency. Moreover, our proposed CP3ER uses no additional losses or exploration strategies. On the right part of Figure 4, we compare the mean, Interquartile Mean (IQM), median, and optimal gap of these methods. CP3ER has significant advantages in all metrics and has more minor variance. This means that CP3ER has better training stability.
Hard-level tasks in DeepMind control suite. We also evaluate CP3ER on 7 challenging tasks [53, 57] in the DeepMind control suite. It should be noted that all tasks here only train 5M frames, rather than the commonly used 30M frames in other work[53, 57]. This means it is a very hard challenge. From the results on the left part of Figure 5, it can be seen that most methods have yet to learn effective policy within 5M frames. Our proposed CP3ER surpasses the performance of all methods without relying on any additional loss or exploration strategies. Moreover, it has significant advantages on all metrics including mean, IQM, median, and optimal gap.
@@ 6.2.1 Can policy regularization improve the behavior of the policy during training?
Action distribution analysis with toy example. In order to further explore the impact of policy regularization on the training, we borrow the 1D continuous bandit problem [9] to analyze the policy behavior. The green curve in Figure 7a (a) shows the reward function. Within a narrow range of actions, the agent receives higher rewards, while within a broader range, the agent can only receive suboptimal rewards. Therefore, the policy needs strong exploration ability to achieve the highest return. We compare Gaussian policy with entropy regularization (MaxEnt GP), Consistency-AC[33] and consistency policy with entropy regularization (MaxEnt CP), and record the action distribution during the training. As shown in Figure 7a, Consistency-AC quickly converges to the local optimal value with the Q-loss. Policy regularization ensures the diversity of action distribution during consistency policy training, preventing the policy from falling into local optima too early. Moreover, consistency policy has robust exploration compared to the Gaussian policy and achieves higher returns.
Dormant ratio analysis. We have shown that the Q-loss can rapidly increase the dormant ratio of the consistency policy network, leading to a loss of policy diversity. In order to analyze whether entropy regularization can alleviate the phenomenon, we record the dormant ratios of the policy networks during the training in 3 tasks. The results are shown in Figure 8. Compared to the rapid increase in the dormant rate in Consistency-AC, CP3ER has a lower dormant ratio, which means that entropy regularization can effectively reduce the dormant ratios of consistency policy.
@@ 7 Conclusion
In this paper, we analyze the problems faced by extending consistency policy to visual RL under the actor-critic framework and discover the phenomenon of the collapse of consistency policy during training under the actor-critic framework by analyzing the dormant ratio of the neural networks. To address this issue, we propose a consistency policy with prioritized proximal experience regularization (CP3ER) that effectively alleviates the training collapse problem of consistency policy. The method is evaluated on 21 visual control tasks and shows significantly better sample efficiency and performance than the current SOTA methods. It is worth mentioning that, to the best of our knowledge, our proposed CP3ER is the first method to apply diffusion/consistency models to visual RL tasks.
@@
8 Acknowledgments
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 62103409, No. 62136008, No. 62293545 and in part by the International Partnership Program of the Chinese Academy of Sciences under Grant 104GJHZ2022013GC.
Appendix A More Results on Dormant Ratios
To demonstrate the phenomenon of policy degradation in Consistency-AC, we conduct experiments with different settings on more tasks and analyze the effect on the dormant ratio. For analyzing the impact of non-stationary data distribution on consistency model training in online RL, we employ 3 classic tasks from D4RL. The results are shown in Figure 10. It can be seen that the non-stationary distribution caused by online training does not significantly affect the dormant ratio of the policy network.
For analyzing the impact of the loss function and high-dimensional state input on the consistency policy, we conduct experiments on 5 tasks separately. Among them, the results of 2 tasks are presented in the main part, and the results of the remaining 3 tasks are shown here.
Figure 11 shows the impact of different loss functions on the performance and the dormant ratio of consistency policy. Compared to SAC policy, consistency policy with Q-loss achieves the higher dormant ratio and the worse performance. Especially on the Finger-turn-hard task, the policy under the Consistency-AC framework hardly learn any meaningful behavior, and the dormant ratios remain at a high value. Figure 12 shows the impact of different observation inputs on Consistency-AC. Compared to state-based settings, image-based policy is more difficult to learn meaningful behavior, and its dormant ratio also maintains a higher value.
Appendix B Implementation Details